### v2

# A Practical Cross-Platform, Multi-Algorithm Study of Quantum Optimisation for Configurational Analysis of Materials

Abstract
Quantum computers show potential for achieving computational advantage over classical computers, with many candidate applications in the domains of chemistry and materials science. We consider the well-studied problem of configurational analysis of materials and, more specifically, finding the lowest energy configuration of defective graphene structures. This problem acts as a test-case which allows us to study various algorithms that are applicable to Quadratic Unconstrained Binary Optimisation (QUBO) problems, which are generally classically intractable exactly. To solve this problem, we implement two methods, the Variational Quantum Eigensolver (VQE) and Quantum Annealing (QA), on commercially-available gate-based and quantum annealing devices that are accessible via Quantum-Computing-as-a-Service (QCaaS) models. To analyse the performance of these algorithms, we use a toolbox of relevant metrics and compare performance against three classical algorithms. We employ quantum methods to solve fully-connected QUBOs of up to variables, and find that algorithm performance beyond this is restricted by device connectivity, noise and classical computation time overheads. The applicability of our approach goes beyond the selected configurational analysis test-case, and we anticipate that our approach will be of use for optimisation problems in general.
1 Introduction
Quantum mechanics provides the framework for understanding chemistry as we know it. The foundations that let us predict the behaviour and properties of any chemical species have been established for almost a hundred years. That the problem size grows exponentially with the number of particles led Dirac to note, ‘The exact application of these laws leads to equations much too complicated to be soluble’, in 1929 [1]. Hence, despite having a framework, it is difficult for classical computers to simulate chemistry efficiently. Quantum computers offer an alternative platform for solving such problems, building from Feynman’s proposal in 1982 of constructing a computer based on quantum hardware [2].
There are several types of chemistry problems that could be considered as candidates to be solved using quantum algorithms, including fermionic simulation [3] and combinatorial optimisation problems. Combinatorial optimisation problems in particular are ubiquitous across science and industry, motivating quantum optimisation as a highly active field of research and development, and a candidate for near-term quantum advantage [4, 5]. In this study we explore the problem of finding the lowest-energy configuration of defective graphene structures, based on the formulation found in Ref. [6]. Suppose we are tasked with removing a specific number of atoms from a graphene sheet of sites. The goal is to find which atoms, when removed, will result in a configuration with the lowest-energy. The configuration of lowest energy will be the one with the maximum number of remaining atom-atom bonds. This is an instance of the densest k-subgraph problem, with many applications beyond configurational analysis [7]. In this work, we consider only the case of removing atoms, and thus .
To solve the problem we first formulate it as a QUBO [8]. QUBOs are paradigmatic formulations capable of encoding a variety of optimisation problems and known to be nondeterministic polynomial time hard (NP-hard) to exactly solve in worst case[9]. Unless PNP, these problems are intractable for classical computers whereas, for quantum computers, modest speed-ups are possible [10, 11], while they remain unlikely to be tractable [12, 13] in the worst case. Nonetheless, it is believed that the correlations arising in real world problems are complex enough to generate hard instances and at the same time provide some structure to be exploited by a quantum computer to potentially solve the problem faster or achieve better solutions (e.g. [14, 15]).
Quantum algorithms for optimisation problems exist for both gate-based quantum computers and quantum annealers. In this work, we apply two well studied practical approaches, variational quantum algorithms (VQAs) and quantum annealing algorithms (QAAs). Despite the simplicity of these methods, and despite efforts in investigating VQAs and QAAs in real quantum hardware, there are significant challenges in using these approaches to achieve quantum advantage in optimisation. These include difficulties such as the approximate nature of these algorithms, issues that inhibit their performance such as barren plateaus [16] and non-adiabatic effects in QAAs [17]. There are also more practical issues around, for example, noise and decoherence [18] and optimisation of controls. As a result, there are many open questions that must be addressed before quantum advantage may be achieved. Valuable insights for dealing with these issues can be attained by experimental implementation of these algorithms on real near-term quantum devices.
Motivated by these challenges, in this work we present a systematic testing of performance of VQA and QAA methods on commercially-available quantum hardware: a gate-based QPU (ibm_fez) and a quantum annealer (D-Wave Advantage). We use the graphene defect problem as our test-case. The simplicity of this model and its use in a previous pedagogical study [6] removes a level of complexity in understanding our analysis, making it more accessible. However, most of our analysis applies to optimisation problems in general.
Our main contributions are:
We synthesize an application-focused benchmarking framework for quantum optimisation algorithms. We analyse several performance metrics, carefully selected to ensure the framework applies to both digital and analog methods. While this framework is built around a specific test-case, the selected metrics remain test-case independent. As such, we anticipate the framework will be useful across various problems and research themes.
We apply this benchmarking framework in practice to the selected test-case and at sizes that reveal the limitations in the implementation of VQAs and QAAs.
Briefly, the assumptions and limitations of our work are as follows: we assume a simple energy model, specifically the one studied in Ref. [6], and therefore our work is not directly applicable to solving real-world problems. We encode the problem constraint into the QUBO formalism using penalty terms, rather than solving a hard-constrained version of the problem. Furthermore, we were limited to using some of the best commercially available quantum devices. We did not apply error mitigation or other techniques that could improve algorithm performance (as listed in Section 6).
In Section 2 we present other recent efforts in benchmarking optimisation algorithms and real world implementations. We then formulate our specific problem in Section 3. In Section 4 we describe our methodology for the cross-platform benchmarking including the chosen metrics, whose results we present in Section 5. Finally, in Section 6 we discuss open questions and future directions. Additional details of our methodology are included in the Appendices.
2 Related Work
Progress has been made in solving quantum chemistry problems with classical hardware, by exploiting domain knowledge about the structure of specific problems [19]. While a standard approach to solving electronic structure problems on classical computers is density functional theory (DFT) [20], we instead use the QUBO formulation of the problem given in Ref. [6]. Some of the best classical methods for solving QUBOs include linear programming (e.g. [21]), greedy algorithms (e.g. [22]) or heuristics (e.g. [23]).
More specifically, Ref. [6] demonstrates how QAA can be applied to solve a version of the graphene defect use-case formulated as a Densest k-Subgraph (DkS) problem (see Appendix A for more detail on its formulation and complexity). This problem is NP-hard for general and bipartite graphs of degree up to [24]. In [6] an atom lattice is encoded as a fully-connected QUBO and executed on quantum annealers. In our work we execute this on larger lattice sizes, using quantum annealers as well as gate-based quantum computers and classical methods. Another related piece of work [25] studies a slightly different formulation of the problem where the goal is to minimise the number of dangling bonds (i.e. connections between vacancies and carbon atoms in the lattice), again executed on a quantum annealer. The same DkS problem as in [6] has also been investigated in a biology-inspired test-case [26] on randomised bipartite graphs, where the topology of the D-Wave device is identified as the main limitation for applying it on quantum hardware.
Ref. [5] provides a comprehensive review of the current state of quantum optimisation, highlighting the challenges and future prospects of the field. The review outlines best practices for making fair comparisons when solving optimisation problems with different methods. The authors showcase recent publications where optimisation problems with varying numbers of qubits and problem connectivity (known as density) are solved using VQAs, comparing each solution’s quality through the approximation ratio metric. The most common problem connectivity involves 3-regular graphs (where each node is connected to three others), which aligns well with quantum hardware layouts. Studies that explore higher qubit counts typically maintain this low problem connectivity. Increasing connectivity for larger qubit systems requires additional SWAP gates, leading to increased hardware noise. However, Refs. [27, 28, 29, 30] all consider Sherrington-Kirkpatrick problems with 100% density, which can be reformulated as QUBO problems. These studies emphasize both the challenge and the importance of choosing standardised metrics to ensure that benchmarks are fair and results are comparable. The chosen set of metrics must fully capture the diverse behaviors of different algorithms. Our work applies the best practices outlined in this review to a particular problem, which stands out in terms of qubit count and high density compared to the other highlighted studies. We achieve a competitive mean approximation ratio with VQE on a QPU relative to the results reported in Table IV of Ref. [5]. While Ref. [5] primarily focuses on gate-based algorithms, we also address questions regarding how QAAs can be fairly compared to variational gate-based methods.
Other studies [31, 32] compare QAAs and gate-based variational methods for specific use-cases. However, a comprehensive study that includes time comparisons, solution quality, and scaling analysis has yet to be conducted. A recent paper [33] compares the Quantum Approximate Optimisation Algorithm (QAOA) to quantum annealing and claims that QAOA, when run on a real gate-based device, can outperform quantum annealing on a 5-regular max-cut graph problem. On the other hand, Ref. [34] argues that these comparisons could be seen as ambiguous, explaining that the reported results from QAOA were enhanced by post-selection. These studies further highlight the need for transparent and fair benchmarking.
3 Problem Formulation
We are focusing on the problem of finding the minimum energy of defective graphene structures. This problem is discussed in more detail in Appendix A. We model the system as an -site hexagonal structure, as shown visually in Figure 1 ( in this example). The first graphene structure explored is a supercell, a unit cell containing two carbon atoms repeated in a arrangement, which corresponds to an variable QUBO. We also investigate other supercells, which translate to QUBOs of varying size. Each (binary) QUBO variable represents a site containing either a carbon atom or a vacancy.
Each of the carbon atoms in Figure 1 can be removed, creating a vacancy that results in the breaking of carbon-carbon bonds. The breaking of these bonds increases the energy of the system’s ground state configuration. The energy increase depends on where the vacancies are introduced. In particular, and following Ref. [6], the energy of the system depends only on the number of remaining carbon-carbon bonds. In this model we impose periodic boundary conditions, making this structure a -regular graph. The supercell’s 2D geometry is maintained for all configurations.
The optimisation problem is, therefore, translated to finding the lowest energy configuration of the system, subject to the constraint of a certain number of vacancies. We work with three vacancies, the highest studied in Ref. [6], treating this number as a constant throughout our work. Holding the number of vacancies constant at three means the space of feasible solutions grows as , where is the number of QUBO variables, which results in polynomial growth. Choosing the number of vacancies to grow as a function of would cause the feasible search space to grow at a larger rate. We experimented with this and found that it made no significant difference, for this particular QUBO instance, to the time scaling of simulated annealing. Therefore, for simplicity we kept the number of vacancies constant at three.
We note that the introduction of vacancies to graphene structures is analogous to doping materials. Hence, this problem may have relevance when optimising configurations in semiconductors, superconductors, thermoelectrics and optoelectronics [35, 36, 37].
To solve our problem on a quantum device, we consider the QUBO cost function:
where is an -component vector containing the binary QUBO variables that represent the different configurations, and is the cost matrix encoding the cost function of the particular problem instance. We aim to solve the minimization problem:
In order to map our problem to the above QUBO, we represent the presence of a vacancy or an atom in each site by being or accordingly. The hexagonal graphene structure (with boundary conditions) can be represented by a graph with adjacency matrix and dimension . Since our objective is to find the maximum number of carbon-carbon bonds left after the atoms are replaced by vacancies, we want to reward (by decreasing the energy of the configuration) only the edges with a in both of their vertices. Therefore, our goal is achieved by choosing a cost matrix that is strictly upper diagonal and contains in its non-zero elements. Parameter is the bond energy, such that adding a single bond lowers the system energy by .
To specify that we must have a certain number of vacancies we can add a constraint term:
to our cost function, where is the number of carbon atoms we want to have in our cell and is a constraint coefficient. Vectors which do not have the number of carbon atoms equal to are penalised in terms of their associated cost value. Bringing all this together, we get:
The -matrix can be read off by comparing Equation 4 with Equation 1. The value of is chosen through a tuning procedure described in Appendix C, where is defined as the QUBO penalty term coefficient. To find the optimal value of the penalty coefficient, we set and search for the optimal for each algorithm. A balance must be struck between enforcing the constraint by increasing and not making too large so that it increases the energy range of the QUBO. Consequently, some solutions can violate the constraint and have an incorrect number of vacancies.
When plotting the problem graph associated with our matrix , we obtain a fully connected graph. This graph is constructed by adding edges between nodes for every . Note that the graph is fully-connected despite the atoms not being bonded in an all-to-all structure; this is because the constraint term in Equation 4 adds in extra on-site and coupling energy terms. To solve the problem on quantum hardware, the graph must be embedded on physical qubits using classical techniques, as described in Appendices D and E. This is one of the major challenges in general when solving combinatorial optimisation problems using quantum computing approaches.
4 Methodology for Multi-Algorithm Benchmarking
Having specified our problem, we discuss the methods and benchmarks we employ. We consider three classical techniques for solving our optimisation problem: brute force, simulated annealing and random sampling. On the quantum side, we use two algorithms: VQE, executed both on a state vector solver and a quantum processing unit (QPU), and quantum annealing implemented on a QPU. Figure 2 provides a schematic overview of the key principles underlying each method. To understand the methodology and solutions from different hardware platforms and algorithms it is important to establish performance metrics which fairly reflect the capabilities of each device and algorithm and that are device and algorithm agnostic. Firstly, the differences between each method are highlighted, along with their limitations. Then, the role of post-selection is discussed and, finally, we outline the chosen metrics and motivation for their use.
4.1 Algorithms and Parameters
Classical brute force simply involves generating all possible solutions to the QUBO problem, calculating their associated energy, and then selecting the optimal ones. Note that we could ‘hard-encode’ the constraint on the total number of vacancies in our calculation by restricting the search space to those configurations that respect the constraint. However, we focus on developing a general framework for the multi-algorithm benchmarking of QUBO optimisation, where the constraint is incorporated as a penalty term as described above. Therefore, in order to ensure a fair comparison, we treat all methods in the same way and do not employ extra steps to enforce the constraint. Brute force is an exact method, so we know the minimum energy returned will be 100% accurate for our model. The major drawback of brute force is the time it takes to exhaustively produce all possible solutions, which grows exponentially with the size of the configuration space we are searching (e.g.[38]).
Simulated annealing [23] is a probabilistic method which draws analogies from the heating and intentional slow cooling of physical systems, such as metals, to avoid structural defects. The cost function is used to compare the current solution against a newly selected one, and improvements on the former are selected. Solutions that do not improve on the current cost function evaluation are also selected with a probability determined by a temperature parameter as a way of combating getting stuck in local sub-optimal minima. The temperature parameter is set to gradually tend to zero, which means fewer worse evaluations will be accepted over time [39].
Many of the methods that we employ here involve hyperparameters whose values must be set. Our hyperparameter search finds the set of parameters for each method that return the highest probability of sampling the ground state. A hyperparameter search could be performed aiming to optimise performance with respect to another metric such as runtime. Our hyperparameter optimisation is a grid-based search, conducted by defining feasible search spaces and checking different combinations of values. The results of the hyperparameter search for all methods in this section are provided in Appendix B.
Simulated annealing provides a probabilistic final solution with a number of samples equal to the number of times it is repeated. Between repetitions only the random seed changes which affects the starting state. Limitations of simulated annealing include the fact that the temperature parameter range and the rate of its decrease must be carefully set. If the rate of decrease is too fast, the algorithm may get stuck in local minima; if the rate of decrease is too slow, the algorithm could take too long to converge due to bouncing around the energy landscape. The simulated annealing hyperparameters we optimise are: (i) the temperature range and (ii) the number of iterations (sweeps).
The final classical method we consider is random sampling. It randomly selects a particular binary vector out of all possible bit strings of bits (like in brute force, we do not exclude from our search space the solutions that have an incorrect number of vacancies) and evaluates its energy with respect to the cost function. This means that each binary vector has probability of being sampled. This is, of course, not uniform with respect to energy as multiple bitstrings can have the same energy. Random sampling provides a useful baseline for solving QUBOs with probabilistic methods, as it uses no technique that exploits the problem structure to improve the probability of finding good solutions.
Turning to our quantum methods, VQE prepares a parameterised trial wavefunction, or ansatz, , where is a set of parameters that can be tuned by changing circuit parameters, to approximate the ground state of a quantum system. The quantum system’s Hamiltonian, , is set to correspond to the classical cost function Equation 1 upon mapping of the binary variables to Ising spins and subsequent quantisation. Measurements are taken over multiple executions of the same circuit, to estimate the energy expectation value. Each execution of the same circuit and subsequent measurement we call a shot. Motivated by the variational principle, a classical optimiser then adjusts circuit parameters to minimize the energy expectation value, , of the Hamiltonian with respect to the variational quantum state .
Each cycle of updating the circuit parameters and estimating energy expectation is called an iteration. The algorithm keeps iterating this process until either the convergence criteria are met based on a tolerance parameter or the maximum number of iterations is reached. This approach of using a quantum resource to prepare and manipulate a quantum state within a classical loop is known as hybrid quantum computing. After convergence, the final circuit can be sampled (again for the same number of shots as each previous iteration) to obtain a probability distribution over the classical bitstring representations [40]. The number of solutions obtained with VQE is equal to the number of shots, which is set to ensure shot noise remains negligible. The whole process described here constitutes an experiment. As in our other methods, there are a number of hyperparameters whose values must be chosen. For VQE, these are (i) the choice of ansatz, (ii) the tolerance of the classical optimiser, (iii) the maximum number of iterations for the classical optimiser, and (iv) the CVaR parameter (described shortly). Note that in this work, we use both IBM quantum hardware and a noiseless classical simulator.
Some of the known limitations of VQE (and VQAs in general) are their iterative nature and their execution time [41]. The iterative nature means that the classical (optimiser) part of the algorithm is susceptible to getting stuck in local minima and not converging on the global minima. This susceptibility is heightened when we consider VQE in the presence of noise and when considering systems of many qubits. We know that the addition of noise can cause the variance of the cost function to decrease [42] which means it becomes harder for optimal parameters to be found. Ref. [43] discusses vanishing variance in loss functions and defines barren plateaus, where the variance vanishes exponentially with the system size. Due to the limited connectivity of current gate-based hardware, many SWAP gates may be needed for QUBO problems with high density. This can result in long circuits, the execution of which is limited by the gate times of the specific device. Long circuits also cause the effects of noise and decoherence to accumulate. This issue is particularly problematic for QAOA which is further described in Appendix D.2 and Ref. [28].
We use VQE with the conditional value at a risk (CVaR) [44] as the objective function of the classical optimiser. In this case, instead of using the usual expectation value as the objective function, which is the average energy of all the measured bit strings, only a fraction of the best bit strings are used to calculate the average energy. This fraction is determined by the parameter , a hyperparameter we tune in our experiments, with corresponding to standard VQE without CVaR. Using CVaR can help convergence on the optimal parameters in fewer iterations [44, 45].
The convergence criteria, set for VQE with the COBYLA optimiser (from the SciPy library [46]), aim for objective values to stop varying within a certain tolerance and within a maximum number of iterations dependent on the problem size. The COBYLA tolerance in the SciPy library is defined as a lower bound on the size of the trust region [47]. We note here that setting stricter tolerance or convergence criteria could improve the VQE’s solution quality in exchange with sacrificing its total runtime. A tolerance informed through hyperparameter search was found to consistently allow VQE and COBYLA to meet the convergence criteria before the maximum iterations were achieved. The number of maximum iterations must be selected based on the time feasibility of running many experiments. We show our convergence plots in Appendix D.
Finally, quantum annealing generates a solution to the given problem by adiabatically evolving a given Hamiltonian to the quantum (target) Hamiltonian that encodes the problem, in an attempt to reach the ground state of the system. Under the adiabatic theorem, if the system starts in the ground state of the initial Hamiltonian (which is an easy to prepare state), it will track the instantaneous ground state if the Hamiltonian varies sufficiently slowly with respect to other energy scales [48]. The duration of the annealing (typically in the range of nanoseconds to microseconds on a real QPU) is known as the annealing time. Longer annealing times increase the likelihood the state obeys the adiabatic principle but also worsens the effect of decoherence. Each quantum annealing device has a range of annealing times that can be set, which is a hyperparameter we optimise in our experiments. The final probability distribution is then sampled by making measurements of the final state. The number of anneals we apply and consequent measurements on the final state we call the number of shots in quantum annealing, as with VQE. Each shot produces a unique solution. The whole process we call an experiment, again similarly to VQE above. Both VQE and QA may sometimes prepare a sub-optimal final state, and we hence repeat experiments for these methods, as we discuss further below.
Due to the limited connectivity of real quantum annealing QPUs, which are usually based on superconducting circuits, variables must be mapped to multiple qubits, forming qubit ‘chains’. A chain is considered broken if the value of one of the qubits in the chain is inconsistent with the others (they should all be the same value as they represent the same variable). Chains are formed via ferromagnetic coupling among qubits, which requires a balance to be struck between increasing the chain strength, so that the chains are not broken, and having the coupling weak enough so that it does not influence the solutions found in the problem [49].
The use of qubit chains in current non-fully connected architectures increases the number of qubits used for each variable of the problem by a quadratic factor, as compared to mapping single qubits to variables. This means that using quantum annealers currently available on the market either restricts us to solve problems that are not fully connected, or, as it is in our case, risks poorer performance when we go to larger problem sizes [50]. The quantum annealing hyperparameters we optimise are: (i) the chain strength and (ii) the annealing time.
4.2 Post-Selection
The output of each algorithm described in the previous subsection is a collection of samples from an underlying probability distribution, which represent possible solutions to the QUBO problem, except for the brute force approach that produces a deterministic outcome. Notice that the QUBO problem is unconstrained and only energetically penalises the solutions that do not adhere to the constraint encoded in its cost function. We can filter out the solutions that do not satisfy the constraint, i.e. the solutions with the wrong number of vacancies, by classical post-processing.
A natural question that arises is: why were these constraint violating configurations sampled at all? Should we not have a large enough penalty coefficient, , so that these configurations are never sampled? The answer is as discussed before - finding the optimal is a balance between making sure these solutions are never sampled and trying to not increase the QUBO energy range too much by penalising the cost of these constraint violation configurations. Often the latter is preferred. Removal of these configurations is achievable by sorting through the bit string solutions and counting the number of present vacancies (represented by zeros). If an incorrect number of vacancies is present in the solution, then the solution is removed. For all the explored methods, the post-selection improves the probability of finding the optimal solution for the original configurational analysis problem.
To perform post-selection for our use-case with a single constraint, a constant number of bitstrings must be checked, equal to the number of samples (shots), which have entries each. Therefore, the scaling of this method is linear in . The time required for post-selection at the QUBO problem size, was found to be approximately for samples, which is the maximum number of samples produced as output in any of our experiments. However, for larger problems, this number of samples may need to scale with to achieve non-zero probability of the optimal solution. As increases, this probability decreases due to increased noise from using more qubits, which we have confirmed experimentally.
The ultimate output from all of these methods is a probability distribution over classical bitstrings or, equivalently, over the configuration energies we sample from, as described above. The classical brute force approach produces the exact solution, but we can associate this with a delta function probability distribution, peaked at the optimal solution. The properties of the full distributions are revealing, and we analyse several of these shortly. Furthermore, this unified output conveniently allows various metrics or summary statistics to be defined that are applicable for all methods. These metrics are important for quick comparisons between different methods and for summarising the important behaviours. We next discuss the metrics considered in this work, which we choose to quantify the performance in terms of both the time to find the solution and the quality of the solution found.
4.3 Performance Metrics
Defining good benchmarking strategies for measuring performance of optimisation algorithms has been extensively studied in both quantum and classical computing contexts [51]. In general, these strategies fall under two distinct categories. The first category involves allocating a fixed amount of resources, such as time or energy, and measuring the quality of solution obtained, using a well-defined metric for it. The second involves measuring the amount of a resource, using an appropriate metric, that is necessary to acquire a solution that achieves a pre-specified threshold (or target) for the quality of solution.
It is therefore clear from the above that, in order to proceed, we need to define metrics of two different types: ones that measure the quality of solution, and metrics that measure the amount of the resource devoted to obtaining a solution. In our case, we select optimal solution probability and approximation ratio as metrics for the quality of solution, and user runtime and QPU time for measuring the resources to solution.
Optimal Solution Probability, – The proportion of times, , the ground state occurs in all solutions:
By ‘solutions’ in the above definition we mean a collection of output samples either before or after post-selection, and this will be specified when the metric is used. This metric is independent of the detailed structure of the probability distribution and describes only the likelihood of successfully finding the solution to the problem.
Approximation Ratio, – The approximation ratio, is defined as:
where is the energy expectation value of the solutions, and are the theoretically calculated minimum and maximum energy values of the cost function over all bit strings that satisfy the constraints, as in Ref. [52].
In our case we calculate AR using the solutions after post-selection. The reason we do this is because before post-selection we can have energies outside the range . These and values are calculated using constrained brute-force techniques. The AR places the average energy found within the context of the whole energy landscape of the problem. We note that there is some variation in the name and precise definition used in the literature for this metric. When exploring new large problems, the theoretical minimum and maximum values of energy may not be readily available without having to solve using an exact method. An option for larger problems is that, in place of the AR, the values of the the average energy and the minimum energy found experimentally by the particular method can be used individually as metrics.
User Runtime – The time experienced by the user when running the experiment:
where is the classical-to-quantum encoding time, is the job processing time (the time it takes to send the job to the desired device, also known as the latency) and the device runtime, which is the CPU or QPU time (or both for VQE on the QPU). In the case of VQE, includes transpilation of the circuit to native gates and optimisation of the physical circuit, whereas, for quantum annealing, is just the embedding time. Note that we do not include the time spent in device queues, and that our separation of quantum and classical time is consistent with Ref. [53].
For classical brute force, random sampling, simulated annealing and VQE with the state vector solver, . For classical brute force, random sampling and simulated annealing is also equal to zero as code is run locally on HPC, such that there is no network latency. The post-selection time, , is another interesting time metric to show for full visibility. As our post-selection time is approximately equal for each method we do not include it in the user runtime, as discussed in the previous sub-section.
QPU Time – The time required on the QPU to reach a solution to the problem:
where is the QPU runtime, specific to running on real quantum hardware.
For D-Wave devices this is the ‘QPU access time’, which includes the QPU programming time [54]. Similarly, for IBM devices this is defined as the time a QPU is committed to complete a job [55], which includes compiling time (converting the transpiled quantum circuit into pulse sequences). Our reported QPU time for the IBM device is calculated by summing the ‘quantum seconds’ of each ‘job’ (in our case a job corresponds to a single iteration of VQE). An alternative method for calculating the QPU is described in Ref. [53]. This metric is important for hybrid algorithms as it separates into CPU and QPU time.
Having discussed the selected metrics, we also note a number of other metrics that are used in the wider literature. Ref. [32], considers the number of violated constraints. For our QUBO there is only one constraint specified - the number of vacancies. The number of times it is violated could be reflected in a ‘validity’ metric: if many solutions violate the constraint then the validity returned will be low. We instead choose to encode information about constraint violation by post-selecting the data to remove solutions that do not obey the constraint, as described above. A metric specific to the quality of the quantum annealing process in current devices is the percentage of broken chains. We report this value in Appendix E, Figure 19(d).
Another metric often used is the time-to-solution (TTS), where this is the time taken for a method to find the optimal solution once with a desired accuracy (often ). TTS, as defined in Ref. [56], is:
Where is the desired accuracy (). The TTS can be calculated from our metrics realising is our user runtime and .
Ref. [57] uses for approximate optimisation, where is the time to reach an energy within a fraction of the ground state energy. is a useful metric for large problem sizes where the ground state may not be found.
Finally, we report error bars on all of our metrics to enable proper statistical interpretation. We have two distinct sources of error in our methods. For our optimal solution probability metric, we have the standard error, SE, which takes into account shot noise when sampling from the output distribution:
Here we create our collection of solutions by aggregating shots across many experiments. This means that the total number of solutions for quantum annealing and VQE is equal to , with being the number of times the experiment is repeated and the number of shots per experiment. While in simulated annealing and random sampling it is equal to the number of times we run the algorithm. For the definitions of shots and experiments see Section 4.1. This is necessary with quantum annealing as the D-Wave devices have a maximum QPU access time, and in order to increase the number of shots to a sufficiently large number, one needs to run the experiment again[58].
We also compute the standard deviation, :
where is a value obtained from the experiment and is the experiment mean. The resulting error bar takes into account errors due to imperfect preparation of the ground state. We therefore repeat ‘experiments’: the entire process of state preparation and generating samples several times (). For each experiment, we compute the values of our quality of solution metrics, and then report the mean and standard deviation as above to capture this variation in state preparation. Note that we also use for our time metrics. Finally, we note that we could simply aggregate all shots across all experiments into one big distribution and not discuss two sources of error. We choose to separate out the two error sources to achieve more detailed insight into these methods, and to allow us to meaningfully characterise the runtime. As is the more dominant error source (notice SE will be small as is large) we report this as our error bars in the main text and report the SE in Appendix C, Table 2.
5 Results
We first solve the variable QUBO and analyse our methods’ performance using the metrics defined in Section 4.3. Figure 3 shows the energy probability distributions obtained by each method before post-selection.
For VQE on the state vector simulator and quantum annealing, the experiment was carried out times, the output samples were accumulated and the distributions renormalised. For VQE on the IBM QPU the experiment was carried out times due to time constraints. We report our metrics averaged across repeat experiments and their error bars as described in the previous section. Finally, as brute force is an exact method, we take it to have idealised values of our quality of solution metrics, a value of for and AR. Figure 4 shows the post-selected distributions, in which solutions that violate the constraint on the number of vacancies are excluded. We see that the distributions have more weight over a range of energies close to the minimum.
Method | Post-Selection | AR Post-Selection | User Runtime (s) | QPU Time (s) | |
Brute Force | N/A | N/A | N/A | ||
Simulated Annealing | N/A | ||||
VQE State Vector | N/A | ||||
VQE QPU | |||||
Quantum Annealing | |||||
Random Sampling | N/A |
To produce the quantum annealing results presented, D-Wave’s physical qubit Advantage System 6.4 was used. The VQE results were produced using the local noiseless IBM state vector solver and the real qubit device ibm_fez, accessible via the cloud from the IBM Quantum Platform. Not all of the physical qubits from the quantum annealing device and gate-based QPU are used: for further details of the qubit embedding procedure see Appendices D and E. The classical brute force, simulated annealing and random sampling were performed on a machine with 32 CPUs and RAM. However, multi-threading was not configured which would better utilize the machine’s capabilities. Acceleration through the use of GPUs would also likely have been beneficial.
We now discuss the results for each method and how they scale for larger sizes of the problem in detail. In the following sub-sections we first focus on the results of quantum annealing (Section 5.1), and then the results of VQE (Section 5.2). Classical results are discussed throughout both sections.
5.1 Quantum Annealing Results
The quantum annealing distribution in Figure 3 is skewed closer to the optimal solution than random sampling, so it has better general performance in the sense of solution quality, as reflected in the optimal solution probability metrics in Table 1. The optimal solution, for the rest of this section, being the optimal energy among the solutions that respect the fixed vacancy constraint in the QUBO. However, solutions below the minimum constrained problem energy are also found. This is a consequence of the small penalty coefficient which was found from hyperparameter optimisation. If is large this increases the QUBO energy range and results in a smaller spectral gap in the annealing Hamiltonian. A smaller spectral gap means that more non-adiabatic effects are likely, which can cause erroneous results.
Comparing the quantum annealing distribution in Figure 3 with the post-selected distribution in Figure 4, we see that the post-selected distribution has support over a smaller range of energies because high-energy/less-optimal solutions are invalid and are removed, as expected. The post-selected distribution for quantum annealing is closer to the optimal solution than random sampling but further from it than simulated annealing. This distribution includes energies that are not optimal, which is reflected in the AR metric.
In Table 1 it can be seen that the QPU time for quantum annealing only makes up a fraction of the user runtime. In Appendix E we dissect the user runtime and show that the majority of this is coming from the embedding procedure (Figure 19(a)). This motivates us to explore more efficient embedding techniques, which we do in the next section as well as solving larger sizes of QUBO.
5.1.1 Scaling to larger problem sizes: Quantum Annealing
We now discuss how quantum annealing runtime scales as a function of problem size. The time spent by quantum annealing on each problem instance is partly determined by the number of anneals (shots) and the annealing time. The number of anneals is chosen so that our results are statistically significant, and the annealing time is found through hyperparameter optimisation. Figure 5 displays the user runtime for simulated annealing and quantum annealing (with two different embedding techniques) as a function of the number of QUBO variables, . The quantum annealing embedding technique used for the results in the previous section was minor embedding, D-Wave’s default mapping scheme, which is a heuristic method. An alternative embedding technique, ‘clique embedding’ [59], can be used for QUBO problems where the QUBO graph forms a ‘clique’, which means that it has full connectivity (100% density). Clique embedding is an algorithm that runs in polynomial time in the worst case[60, 61, 59], which has uniform chain lengths (or close to uniform) [60]. We observe near-constant scaling with clique embedding in Figure 19(a). At larger problem sizes this may start to grow polynomially.
As discussed previously, the primary contributor to the user runtime of quantum annealing using minor embedding is the embedding time, detailed in Appendix E and explicitly shown in Figure 19(a). As problem size increases, the required average qubit chain length also increases, shown in Figure 19(c). Consequently, the heuristic minor embedding technique must explore more mapping possibilities, leading to poor scaling. This can also explain the large standard deviation error bars observed in Figure 5: there is more variability possible when the minor heuristic embeds larger problems and requires longer qubit chains. A new embedding is found each time the minor heuristic is performed, which can give variable performance in terms of time and quality of solution. In future work, the minor embeddings which yield good performance could be saved and re-used for multiple experiments.
While considering the runtime scaling, it is important to not lose sight of solution quality. Quantum annealing with minor embedding was able to solve up to the 72 variable QUBO while still returning some non-zero . On the other hand, clique embedding was able to solve up to the 50 variable problem while returning non-zero . We include results with as reduced opacity data points in Figure 5 to give insight into how quantum annealing scales for larger problems. The quality of solution and time metric results are shown in Table 5, where we see that, in general, minor embedding returns a higher but with very large standard deviation errors.
Turning now to the scaling behaviour of classical simulated annealing, we solve the QUBO with up to 338 variables. None of our other methods are performant for such large problems, so we verify these simulated annealing solutions with a modified version of brute force which exploits the structure of the problem. Instead of checking an exponentially-large number of states with brute force, for our particular problem we can search through just the states with the specified number of vacancies. For example, this means that for our 32 variable QUBO where we are looking for solutions with three vacancies that this version of brute force now only has to search through configurations.
We note that we only use this constrained search space for brute force to verify these larger problem sizes. It is fairer practice to compare the brute force method that searches through all configurations with our other methods as this is the space they are also searching. Using this modified constrained version of brute force effectively changes our entire problem from QUBO to a hard-constrained DkS optimisation problem.
We could equally reduce the search space for our variational quantum methods in the constrained case using a space-restricting ansatz [62, 63, 64] or quantum Zeno dynamics [65]. We could also consider constrained quantum annealing [66]. Another strategy to reduce the search space is to consider symmetrically equivalent structures only once. With VQAs, for example, this could be done with a symmetry respecting gate-set [67].
Returning to our results, we observe polynomial time scaling for simulated annealing. This is not surprising as NP-hard problems have subsets of instances that can be solved in polynomial-time by heuristic algorithms [68, 69]. These results highlight the importance of selecting QUBOs which are hard to solve using common classical methods in order to fully stretch quantum computing methods.
For quantum annealing, the theoretical time scaling is bounded by the adiabatic limit, which depends on how the energy gap between the ground state and first excited state of the annealing Hamiltonian closes with system size. This places upper limits on the computational complexity of problems that can be addressed. For first order quantum phase transitions, the annealing rate is bound exponentially with system size. For second order quantum phase transitions, the annealing rate is instead bound polynomially with system size [70]. Therefore, the best scaling we can expect from quantum annealing is polynomial time, which we see experimental evidence of with clique embedding.
5.2 VQE Results
Turning now to the other quantum method we considered, we see from the distributions in Figure 3 that both VQE on the state vector solver and on the QPU have small probability of sampling large energies, although the distributions do skew ‘to the left’ compared to naive random sampling.
When comparing VQE on the QPU to the state vector solver we can see that the state vector solver has higher . This is due to VQE on the QPU sampling higher energies more often, which is likely a result of noise. Otherwise, their distributions match well. Noise can alter the shape of the cost function landscape, introducing false minima or saddle points. The classical optimiser for VQE on the QPU could be getting stuck in these local minima, failing to converge on the global minima exactly. Post-selection improves for both the state vector solver and the QPU through removal of these higher energies.
A feature of the results in Table 1 is that VQE has a large standard deviation error, , associated with the metric. As discussed, captures the inconsistency in convergence with the classical optimiser in VQE. Convergence heavily depends on the initial starting point of the circuit parameters, which is randomly set for each repetition. Sometimes these starting parameters are unfavourable, perhaps close to local minima, and result in VQE struggling to find the global minima.
Inspection of the time metrics in Table 1 reveals that the user runtime for VQE state vector simulation is more than times longer than that of simulated annealing. VQE on the QPU has a user runtime which is over 2 orders of magnitude larger than the state vector simulation. Only of this time is made up of QPU time. The rest of this user runtime is coming from classical processes, such as: compilation and classical optimiser time.
Strategies to reduce this user runtime could include using a less-optimised transpilation search at the expense of solution quality. Another recent technique introduced by IBM is fractional gates [71]. Fractional gates are essentially new native gates which remove the need for and rotation gates to be decomposed into many native gates. This can reduce circuit depth significantly. Our initial experimentation revealed that they can reduce the user runtime by up to a factor of two. Another technique that can be used to reduce gate depth is AI transpilation [72]. As discussed in Section 4, less strict convergence criteria would result in less iterations (and shorter user runtime), again sacrificing solution quality.
As we did for QA, we now explore solving different sizes of our QUBO problem with VQE on the QPU.
5.2.1 Scaling to larger problem sizes: VQE QPU
Performing a scaling analysis for VQE on a real QPU proved challenging, primarily due to runtime increases as the problem size is scaled up. Beyond 50 variables, VQE failed to return non-zero . However, at the 72 variable problem size, convergence criteria were met and sub-optimal solutions were found (distribution and convergence plots are shown in Figures 13, 14, 15, 16 and 17, as well as the performance metrics for the different problem sizes in Table 4). At large problem sizes, more qubits are used and errors accumulate causing poor convergence to a state which has no overlap with the ground state. The extended runtime relative to the allocated device access time also meant that multiple experimental runs were not possible. Moreover, proper hyperparameter optimisation was not possible either due to this runtime, meaning that hyperparameters either need to be imported from smaller problem instances ran on the state vector solver or extrapolated from them.
Due to these complications, we do not report a practical scaling analysis for this problem solved with VQE on quantum hardware. However, we can make some theoretical comments on the expected scaling behaviour of VQE and VQAs in general. Notably, VQE is a heuristic algorithm, and to our knowledge, there is a lack of evidence of provable speed-ups for VQAs on optimisation problems. Ref. [73] does claim a provable speed-up for QAOA applied to a specific symmetric optimisation problem and Ref. [14] also provides empirical evidence that QAOA offers a modest polynomial speedup over leading classical solvers for random k-SAT, which is a satisfaction problem rather than an optimisation problem (the authors of Ref. [74] believe their results on the equivalence of QAOA to quantum annealing apply for this k-SAT instance, which implies that quantum annealing should at least match QAOA’s performance in terms of AR). Another quantum algorithm which has provable speed-ups for optimisation problems is decoded quantum interferometry (DQI) [15], which contains the quantum Fourier transform as a sub-routine and is therefore a fault tolerant algorithm.
Regarding near-term algorithms, it remains to either find structured instances of problems where provable speed-ups exist or to experimentally find possible modest speed-ups. It is also worth noting that using a one-qubit-per-variable encoding makes it impossible to scale to large problems with current hardware due to qubit count limitations. Efforts in multi-variable to qubit encoding are investigated in Refs. [75, 76]. In Appendix D.1 we discuss barren plateaus and classical simulability in the context of our ansatz.
6 Conclusions and Outlook
We introduced a systematic approach for analysing the performance of various algorithms for solving QUBO problems. This approach is applicable to other use-cases and problems in general. We applied it to the configurational analysis problem, determining the energy of defective graphene structures. We defined performance metrics, considering both solution quality and runtime, that are hardware and algorithm agnostic. We then explored the scaling of the problem to larger sizes. Our results show that simulated annealing was the best-performing method in both solution quality and runtime, with a polynomial scaling in time. It can also access large problems of hundreds of variables, in contrast to the quantum methods. We also find evidence of polynomial scaling with quantum annealing, solving QUBOs up to variables. Of the two embedding techniques that we tested, clique embedding was found to be optimal in terms of runtime.
We encountered challenges in scaling the VQE methods for several reasons, namely lack of convergence within the maximum number of iterations, convergence to states with no overlap with the ground state and limited experimental repetitions. However, we were able to solve a dense QUBO with 50 variables and obtain sub-optimal solutions for a 72 variable dense QUBO. A myriad of techniques that could further improve the performance of gate-based VQAs exist, such as:
Warm start [77], where instead of choosing random initial parameters for the ansatz, the starting guess comes from a classical solution to a relaxed (continuous variable) version of the problem.
Concentration of parameters (also referred to as transfer of parameters) [78], where, similar to warm start, the starting parameters of the VQA are chosen based off information from other, easier to solve, problems.
Ascending-CVaR objective functions can also be used, where the CVaR parameter is increased in each iteration [79].
The use of a Gibbs objective function [80] follows similar logic to that of CVaR, where the objective function is formed using the Gibbs exponent.
Recursive QAOA [81], which iteratively reduces the problem size by fixing variables.
Techniques also exist for quantum annealing, such as reverse annealing [85], using different annealing schedules and the variational adiabatic algorithm [86]. The effect of applying error mitigation techniques on both platforms could also be explored further. However, implementing all of these techniques to see improvements in performance was beyond the scope of this work and we leave this as a future research direction.
We also note that simulation of quantum annealing is possible using path integral Monte Carlo (PIMC) methods [87, 88], which may be of use for verifying small-scale quantum annealing solutions in future studies.
Ideally, we want to find instances of QUBO that are hard for classical techniques but suited to quantum algorithms and hardware. What makes classes of QUBO hard for algorithms in general is an open question. Certainly, solution degeneracy, which dictates the number of local minima in the cost function landscape is a contributing factor [89, 90]. Through our study, we can make some observations about what makes QUBOs hard to solve on quantum devices: QUBO density, which relates to device connectivity and, for quantum annealing, the total energy range of the problem which corresponds to preferred penalty terms and limits the annealing rate. Other fertile areas for future work could focus on constrained optimisation [62, 63], optimisation problems with symmetries [73] and multi-variable to qubit encoding [75, 76].
Quantum hardware is constantly improving, with many companies producing ambitious roadmaps promising more and high-quality physical qubits, and early steps towards quantum error correction. We leave finding potential speed-ups to future experimental works and hope that they can follow our framework for analysing the performance of different methods when solving optimisation problems.
Finally, we would like to stress that this paper is not intended to be a complete comparison of gate-based devices and quantum annealing. It is rather a framework for the comparison of ‘end-to-end’ approaches to solving optimisation problems. Our framework compares whole approaches to solving a problem, including the algorithm and hardware platform, and does not focus on the hardware alone.
Acknowledgements
The authors would like to thank Bruno Camino, Stefan Woerner, Andrew King and Robert Cumming for useful discussions, and Vincent Graves for proof-reading the manuscript.
Code Availability
The code used to generate the experimental results presented can be found in the public GitLab repository [91].
Author Contributions
Phalgun Lolur conceived and planned the project. Kieran McDowall (KM) carried out the investigation. KM and Theodoros Kapourniotis analysed and interpreted the results with support from Chris Oliver. Konstantinos Georgopoulos was responsible for overall supervision of the work. KM prepared the manuscript which was then proof-read and revised by all authors.
Appendix A Graphene Defect Problem and its Computational Complexity
Our configurational analysis use-case is finding the energies of defective graphene structures. More specifically, given a graphene sheet which we represent by a hexagonal carbon lattice of atoms with boundary conditions, we remove some of the carbon atoms, creating vacancies on the corresponding sites of the lattice. This results in breaking of carbon-carbon bonds, or simply ‘bonds’ hereafter, between the atom we removed and its neighbours in the lattice. The energy of the resulting structure can be determined by the atoms left and their connectivity. Different approaches in determining this energy function and thus the corresponding optimisation problem have been taken in [25] and in [6].
We use the same approach as in [6]. Here the problem is the following: If we are asked to remove a specific number of atoms (i.e. create a fixed number of vacancies), the goal is to find which atoms, when removed, cause the maximum number of bonds remaining in the structure. We can model the above as a graph theoretical problem on a graph , with . Each vertex represents a site, which can contain an atom, in which case , or a vacancy, in which case , where . The edges represent the hexagonal carbon lattice of graphene. The condition of our problem is that the size of , i.e. the number of vacancies, needs to be fixed, . Then the problem is to determine the set , i.e. the position of the vacancies, so that the bonds, which are the edges in that connect the vertices of , are maximised.
This a very common problem in complexity theory called the Densest subgraph (DkS) [7]. Notice that the configuration space, i.e. the number of possible selections of the vertices of the set of fixed size out of sites scales with . For constant (or ) the configuration space is polynomial on and therefore exhaustive search gives a polynomial time algorithm (note: exhaustive search for the QUBO formulation of the problem is exponential in time as the search space is ). For arbitrary and for general graphs the problem is known to be NP-Hard, even for graphs of degree as ours [24]. The proof is by simply observing that that the maximum clique problem can be reduced to DkS. The best known polynomial approximation algorithm has approximation ratio [92]. In the special case of bipartite graphs, such as our hexagonal carbon lattice with boundary conditions, it has been proven that the DkS problem remains hard [93]. It is also known that densest problems (more edges in the original graph ) are harder in general [26].
A variety of algorithms to solve this problem, including heuristics, exist, and quantum algorithms have also been used [6, 26]. In [26], based on a biology use-case, randomised bipartite DkS problems with and are tried on Chimera-architecture D-Wave annealers finding that they cannot handle this problem. We use the more connected D-Wave architecture. For an account of classical algorithms solving this problem see the survey in [94]. It has been shown that tabu search heuristics achieve better performance than simulated annealing and greedy randomised search [95]. Also, the variable neighbourhood search heuristic performs better than tabu search for sparse initial graphs, such as ours.
A different approach in modeling the problem is taken in [25], which requires us to double the number of qubits. Here only the so-called ‘dangling bonds’, which an atom is connected to a vacancy, increase the energy of the structure and therefore its instability. Now, if we are asked to remove a specific number of atoms (a condition not enforced by the authors of [25]), the goal is to find which atoms, when removed, cause the minimum number of dangling bonds. The graph theoretical problem is therefore slightly different. Again we have a graph , with with vertices representing sites and the edges the carbon lattice, and again represents the set of vacancies. Again the problem is to find a partition of into two fixed sized sets and , but the goal now is to minimise the number of elements for which and (in other words the edges that connect one member of with one member of , are minimized). This is the well-studied minimum bisection problem (MBP), known to be NP-Hard for general graphs [96, 97]. A typical formulation of the problem is to choose . In the special case of regular graphs, such as our hexagonal carbon lattice with boundary conditions, it has been proven that the MBP problem is hard to even approximate [98]. We think that this is a very promising version of the problem to be considered in future work.
Appendix B Hyperparameter Search
Our hyperparameter search finds the set of parameters for each algorithm which return the highest probability of sampling the ground state. The hyperparameter optimisation is a grid-based search, conducted by defining feasible search spaces and checking different combinations of values. Other, more rigorous strategies for hyperparameter optimisation exist, such as random search and Bayesian optimisation. Use of these methods could yield better hyperparameters. Results are plotted for each algorithm in Figures 6, 7, 8, 9. Hyperparameters for VQE on the QPU were informed by VQE on the state vector solver. Other schemes exist to find the optimal penalty terms. A more systematic approach is explored in Ref. [99] for finding these terms.
Appendix C Technical Information and Additional Data for Simulated Annealing
Method | Post-Selection | AR Post-Selection | User Runtime (s) | QPU Time (s) | |
Brute Force | N/A | N/A | N/A | ||
Simulated Annealing | N/A | ||||
VQE State Vector | N/A | ||||
VQE QPU | |||||
Quantum Annealing | |||||
Random Sampling | N/A |
Method | Post-Selection | AR Post-Selection | User Runtime (s) | QPU Time (s) | |
Brute Force | N/A | N/A | N/A | N/A | N/A |
Simulated Annealing | N/A | ||||
VQE MPS | N/A | ||||
VQE QPU | |||||
Quantum Annealing | |||||
Random Sampling | N/A |
D-wave’s ‘SimulatedAnnealingSampler’ was used, where all hyperparameters found from the search at the 18 variable problem size were also used for the larger problem sizes. For the rate of decrease of the temperature parameter, a default ‘geometric’ annealing schedule was chosen, and a default number of solution updates or iterations (number of sweeps) was found from the hyperparameter search, which was [100]. The optimal penalty term coefficient for simulated annealing was found to be , which was found via hyperparameter optimisation.
Appendix D Technical Information and Additional Data for VQE
QUBO Variables | Post-Selection | User Runtime (minutes) | QPU Time (minutes) | |
The number of shots was set to 10,000 for both the real device and for the state vector solver. To solve the QUBO problem using VQE, each variable is initially mapped to a single physical qubit. Qiskit’s circuit transpilation is used when running on the real device, with the optimization_level=3 [101]. The qubit connectivity plot for ibm_fez is shown in Figure 10(a), where the qubits that were used are highlighted in red. The qubits that were used were chosen based off the error rates of each qubit, which can vary with device calibration. for VQE on the QPU should be minimal as qiskit serverless was used.
The classical optimiser used was COBYLA for the state vector solver and for the real device. For the 18 variable problem an optimiser tolerance of 1 was set as this was the value where successful termination was achieved within the number of specified iterations (250). The chosen tolerance and maximum number of iterations is dependent on the problem size and informed by hyperparameter search.
The circuit ansatz used was Qiskit’s ‘RealAmplitudes’ [102]. For this study, only one alternating layer (repetition) was used, along with the default entanglement pattern ‘reverse_linear’. Measurements are made in the computational basis (single qubit Pauli-Z basis), such that a 0 or 1 is measured depending on the overlap with the Pauli-Z eigenstates and .
[…]
@C=1em @R=1em
[…]
D.1 Ansatz, Lie Algebra, Barren Plateaus and Classical Simulability
The ansatz used was qiskit’s RealAmplitudes circuit, shown in Figure 11, which is commonly used for chemistry applications or classification circuits in machine learning [102]. It is particularly suited for problems with real-valued solutions (since the ansatz only generates real amplitudes, no complex amplitudes). The circuit consists of alternating layers of Y rotations and CX entanglements. Figure 12 shows the ansatz for the 2 qubits.
[…]
The Dynamical Lie Algebra (DLA) for this circuit is given by all possible nested commutators between the generators , until no new independent skew-Hermitian operator is generated. The first two elements of the DLA are and (which we will write as and ), since . By noting the third element of g is . To find other elements of the DLA we now take the commutator of our current elements. We can do this in Pennylane [103] to give the DLA:
Considering the additional two gates produces no new elements in the DLA. Therefore . The dimension of the special unitary group is , which for gives . The Hilbert space is and for full universality, we require the available gates to span all of . That means when we have all unitaries of available to us, we can reach any state in Hilbert space from any other state. Therefore, the RealAmplitudes ansatz is not universal.
The dimension of the DLA for the RealAmplitudes with reverse linear entanglement scales polynomially with n. This can be seen by the fact that each contributes a local generator , therefore, n generators. Each gate introduces a new generator (e.g. ) on the target qubit. The number of these generators scales linearly with n because the entanglement is limited to neighbouring qubits. The new number of generators produced under commutation is restricted by the limited connectivity.
Ref. [104] states that barren plateaus do not exist for QUBO problems due to having 2-local observables. A key idea in Ref. [105] is that if there is an absence of barren plateaus this implies that your circuit can be simulated classically in polynomial time (rather than exponential time using the state vector simulator). The simulation algorithm: G-sim [106], exploits low dimensional lie algebra. In summary, our ansatz likely avoids barren plateaus and is efficiently classical simuable.
D.2 RealAmplitudes vs QAOA Ansatz (Hamiltonian agnostic vs informed)
QAOA is popularly used for optimisation problems as it is Hamiltonian informed. The form of the QAOA circuit ansatz is:
where, is the number of layers. and are variational parameters. is the problem Hamiltonian, which encodes the cost function of the optimisation problem and makes it Hamiltonian informed. is the mixer Hamiltonian, typically chosen as , where is the Pauli-X operator acting on the -th qubit. is the initial state, which is an equal superposition of all computational basis states [107].
When comparing different ansatz in our hyperparameter search we found that the RealAmplitudes ansatz returned higher values of (results in Figure 8(e)). The simple explanation for this is that the RealAmplitudes ansatz has more parameters and is therefore more expressive allowing it to find a state with good overlap with the ground state. But what happens if we increase the depth of QAOA so that it has the same number of parameters as the RealAmplitudes ansatz? At , for the variable problem size, QAOA has just independent parameters, but already has a depth of gates, whereas the RealAmplitudes depth is gates (these are the abstract circuits before transpilation). This is due to the fact that our problem Hamiltonian , is fully connected. Consequently, the QAOA circuit has full connectivity (a fully connected entanglement arrangement), implementation of which on the IBM QPU is possible only with the introduction of SWAP gates. Running and simulating these QAOA circuits at can already take a long amount of time which makes increasing quite infeasible.
However, Ref. [108] shows QAOA can have good performance at on a fully connected SK-problem. This is possible due to the concentration of parameters found with QAOA, where parameters at smaller problem sizes can be extrapolated to larger problem sizes. The power of QAOA is that it can perform well with a relatively small number of parameters at low depth which should allow scaling to larger problem sizes. Future work could further explore the practical implementation of QAOA, using concentration of parameters to scale to large, dense QUBOs.
Appendix E Technical Information and Additional Data for Quantum Annealing
To run a QUBO problem on the quantum annealer, each variable must be encoded to the hardware. Given the limited connectivity of the hardware, multiple physical qubits need to be mapped to a single variable. The mapping procedures used were minor embedding, D-Wave’s default mapping scheme, and clique embedding, intended for fully connected problem graphs. The minor embedding mapping to the D-Wave Advantage_system 6.4 QPU is shown in Figure 10(b). The time taken for this embedding procedure on the 18 variable problem was found to be . The percentage of broken chains for minor embedding on the 18 variable QUBO was %.
Figure 19(a) shows the embedding time required for all problem sizes solved. Figure 19(b) displays the QPU access time that the D-Wave device allocates for solving problem sizes of up to 72 variables, and Figure 19(c) presents the average chain lengths used. Full results for quantum annealing using both minor and clique embedding are shown in Table 5 for the larger problem sizes.
QUBO Variables | Embedding | Post-Selection | Runtime (s) | QPU Time (s) | |
Minor | |||||
Minor | |||||
Minor | |||||
Minor | |||||
Minor | |||||
Minor | |||||
Clique | |||||
Clique | |||||
Clique | |||||
Clique | |||||
Clique | |||||
Clique | |||||
Clique |
QUBO Variables | Embedding | Chain Strength | Annealing time (ns) | |
Minor | ||||
Minor | ||||
Minor | ||||
Minor | ||||
Minor | ||||
Minor | ||||
Clique | ||||
Clique | ||||
Clique | ||||
Clique | ||||
Clique | ||||
Clique | ||||
Clique |